find the foci, length of the major axis, and length of the minor axis for the ellips x^2 + 4y^2 = 16.
Answer by lwsshak3(11628)

find the foci, length of the major axis, and length of the minor axis for the ellipse
x^2 + 4y^2 = 16.
divide by 16
x^2/16 + y^2/4 = 1
This is an equation for an ellipse with horizontal major axis.
Its standard form: (x-h)^2/a^2+(y-k)^2/b^2
For given equation:
center:(0,0)
a^2=16
a=√16=4
length of horizontal major axis=2a=8
b^2=4
b=2
length of minor axis=2b=4
Foci:
c^2=a^2-b^2=16-4=12
c=√12≈3.5
foci: (0±√20,0)=(0±3.5,0)=(-3.5,0) and (3.5,0)